# Learning to See and Act: Task-Aware Virtual View Exploration for Robotic Manipulation

### v4

@@ Abstract
Recent vision-language-action (VLA) models for multi-task robotic manipulation commonly rely on static viewpoints and shared visual encoders, which limit 3D perception and cause task interference, hindering robustness and generalization. In this work, we propose Task-aware Virtual View Exploration (TVVE), a framework designed to overcome these challenges by integrating virtual view exploration with task-specific representation learning. TVVE employs an efficient exploration policy, accelerated by a novel pseudo-environment, to acquire informative views. Furthermore, we introduce a Task-aware Mixture-of-Experts (TaskMoE) visual encoder to disentangle features across different tasks, boosting both representation fidelity and task generalization. By learning to see the world in a task-aware way, TVVE generates more complete and discriminative visual representations, demonstrating significantly enhanced action prediction across a wide array of manipulation challenges. To further validate the robustness and generalization capability of TVVE under out-of-distribution (OOD) settings, we construct a challenging benchmark, RLBench-OG, covering various visual perturbations and camera pose variations. Extensive experiments on RLBench and RLBench-OG show that our TVVE achieves superior performance over state-of-the-art approaches. In real-robot experiments, TVVE demonstrates exceptional performance and generalizes robustly in multiple OOD settings, including visual disturbances and unseen instructions. Visual results and code are provided at: TAVP.
@@ 1 Introduction
A general-purpose robotic system is expected to accurately perform numerous manipulation tasks [liu2025aligning]. With the development of vision-language-action (VLA) models [kimopenvla, rt-1, rt-2, livision, cheang2024gr, bu2024towards, li2025bridgevla, bu2025univla, cen2025worldvla, liu2025hybridvla, team2024octo, lin2025onetwovla] and transformer-based architectures [ACT, rt-1, rt-2, hou2024diffusion, chi2023diffusion], many end-to-end robotic manipulation frameworks (e.g., OpenVLA[kimopenvla], $\pi_{0}$[black2410pi0]) have achieved promising results on complex and fine-grained tasks. Typically, these models rely on observations captured from a single view [kimopenvla, chi2023diffusion, shridhar2022cliport] or a few fixed viewpoints [ke20243ddiffactor, shridhar2023perceiver] to guide action prediction. While this may suffice in simple environments with only a few objects, it becomes problematic in cluttered or dynamic scenes—particularly when objects or the end-effector are repositioned during task execution.
@@
Beyond acquiring complete and informative visual features, a scalable and task-sensitive visual encoder is also critical for accurate perception in multi-task settings. Although prior language-conditioned transformer-based methods, such as RVT [goyal2023rvt] and RVT-2 [goyal2024rvt2], enable multi-task capabilities by attending to instruction tokens, their shared encoders still suffer from task interference, especially when handling visually and semantically diverse tasks (e.g., pick an apple vs. open a drawer). This entanglement limits the generalization and scalability of such models across a broad range of manipulation behaviors.
@@
To comprehensively evaluate the effectiveness of our TVVE in enhancing robotic spatial understanding and action generation performance, as well as its multi-task perception and decision-making capabilities, we conducted rigorous benchmark testing on 18 diverse RLBench [james2020rlbench] simulation tasks covering multiple tasks and scenarios. Our contributions are summarized as follows:
  • Multi-Viewpoint Exploration Policy (MVEP) : The MVEP enables robots to explore optimal virtual observation viewpoints (dynamic multi-view re-rendering), effectively addressing occlusion and insufficient viewpoint problems, and enhancing 3D perception capabilities.
@@ 2.1 Multi-task Learning in Manipulation
The core challenges for general-purpose robots lie in task generalization capabilities and multi-task execution in complex environments. In the past, multi-task learning [deisenroth2014multi, ze2023gnfactor, goyal2024rvt2, ma2024hierarchical] has achieved remarkable progress. It primarily consists of two approaches: Modular solutions decompose skills into reusable basic units (e.g., perception, planning, and control modules) [mao2024robomatrix]. Though interpretable, they face the combinatorial explosion problem—as task complexity increases, the cost of designing interfaces and coordinating between modules grows exponentially, leading to system rigidity. For instance, ManipGen [dalallocal] consumed massive training resources to train over 1,000 task-specific experts. Conversely, end-to-end solutions employ a single dense model to directly map perception to action [team2024octo, ze20243d, wang2024poco, chi2023diffusion, shridhar2022cliport, nair2022learning, haldar2024baku]. While simplifying the pipeline, significant differences in visual feature distributions and motion patterns across tasks (e.g., grasping fragile objects vs. tightening screws) cause the model to struggle with convergence during training due to parameter conflicts. The Mixture of Experts (MoE) [shazeer2017outrageously] addresses the scaling bottleneck of large models through a sparse activation mechanism. SDP [wangsparse] integrated MoE into diffusion policies to address multi-task learning and task transfer. Differently, we propose a task-aware MoE (TaskMoE) to resolve feature conflicts between different robotic manipulation tasks, activating distinct task-specific experts for different tasks, thereby enhancing multi-task execution performance and generalization capabilities.
@@ 2.2 Vision-Language-Action Model
Recently, numerous VLA models [shridhar2022cliport, team2024octo, ze20243d, goyal2024rvt2, zhang2025autoregressive] have been proposed. In contrast to traditional pipelines that rely on explicit pose estimation [fang2020graspnet, deng2020self] and path planning [prianto2020path, zhao2025deep], VLA models directly map linguistic instructions and visual observations to low-level actions, substantially simplifying the control stack. With the advancement of large language models (LLMs) and vision-language models (VLMs), these large models have been applied in robotics [kimopenvla, rt-1, rt-2, livision, cheang2024gr, bu2024towards], leveraging their powerful priors to provide world knowledge for robots, thereby enhancing model generalization and performance. However, most existing VLA models operate on a single or fixed viewpoint, limiting a robot’s ability to form comprehensive 3D spatial understanding and thereby constraining action prediction. To overcome this limitation, we introduce a dynamic multi-view re-rendering perception scheme that enhances spatial perception and scene comprehension, ultimately improving execution robustness in complex environments.
@@ 2.3 Reinforcement Learning in Robotics
Reinforcement learning (RL) has achieved significant accomplishments in robotics [tang2025deep]. However, implementing RL starting from scratch typically necessitates complex reward mechanism and training paradigm design [hu2024flare, khetarpal2022toward, luo2024serl, rajeswaran2017learning, uchendu2023jump]. Recently, several works [guo2025improving, lu2025vla] have demonstrated promising results by using RL to fine-tune pre-trained models. To optimize our MVEP (Multi-Viewpoint Exploration Policy), we designed pseudo-environment interactions and a view-aware reward function, enabling policy learning based on a pre-trained model using offline expert data. Furthermore, to adapt between the MVEP and the action policy, we introduce supervised post-training to coordinate the robot’s ability to “See and Act”, thereby enhancing its task execution performance.
@@ 3.1 Overview
Our proposed Task-aware Virtual View Exploration (TVVE) framework aims to identify optimal viewpoints for accurate and robust robotic manipulation, guided by task-specific visual feature extraction. The full pipeline is illustrated in Fig. 2. TVVE takes as input a language instruction, the current visual observations from RGB-D cameras, and the current gripper state. To enable the model to explore the optimal observations from arbitrary viewpoints, we first reconstruct a 3D point cloud of the scene using the input RGB-D images. To narrow down the search space and enable task-aligned view selection, we leverage the coarse prediction stage from RVT-2 [goyal2024rvt2] to generate an area of interest. However, unlike RVT-2, which extracts visual features for all tasks using a shared Multi-View Transformer (MVT), we integrate a Task-Aware Mixture-of-Experts (TaskMoE) module within the MVT to route instructions to specialized expert encoders. This design enables more precise and task-aligned visual feature extraction, which benefits both viewpoint selection and action prediction. Starting from the identified area of interest, we employ a Multi-Viewpoint Exploration Policy (MVEP) to search for optimal camera poses that maximize the visibility of the target object and the end-effector. The selected viewpoints are then re-rendered into image observations and processed using another TaskMoE-based MVT before being passed to the action prediction model. For action prediction, we also upgrade the autoregressive action policy proposed in ARP [zhang2025autoregressive] with our proposed TaskMoE to attain more task-specific feature extraction and action prediction.
@@ 3.2 TaskMoE
To address the inherent heterogeneity of complex manipulation tasks in multi-task learning, where different tasks often require substantially distinct visual representations and action policies, we introduce a task-aware Mixture-of-Experts module (TaskMoE), as illustrated in Fig. 3. Compared to prior MoE-based approaches [shazeer2017outrageously, wangsparse], our TaskMoE introduces two key innovations.
First, instead of relying solely on task identifiers for expert selection, we incorporate richer instruction- and scene-related cues to guide expert routing more effectively, which is crucial for accurate multi-task robotic manipulation. Specifically, as shown in Fig. 3, we design a cross-modality module that employs a cross-attention mechanism to model the interaction between instruction and visual information. The resulting context-aware features are then fused with the task identifier via a Feature-wise Linear Modulation (FiLM) layer [Film], enabling more adaptive and task-sensitive expert selection.
@@ 3.3 Multi-Viewpoint Exploration Policy
In this work, we represent each camera pose using a look-at model [lookatmodel], which decouples the camera position and orientation via spherical coordinates. Therefore, the explored views are parameterized as 5-dimensional vectors $\mathbf{p}^{i}=(\theta^{i},\phi^{i},r^{i},\theta_{\text{up}}^{i},\phi_{\text{up}}^{i})\in\mathbb{R}^{5}$, where $(\theta,\phi,r)$ specify the camera center in spherical coordinates relative to the origin, $(\theta_{\text{up}},\phi_{\text{up}})$ define the orientation of the up vector, and $i\in[0,K-1]$ means the $i_{th}$ view. So, all camera poses are $\mathbf{p}=\{\mathbf{p}^{i}|i=0,...,K-1\}$. Additionally, the camera is perpetually oriented toward the coordinate origin $O(0,0,0)$.
@@ 3.4 Training Strategy
Stage 1: In the first stage, we train a fixed-view variant of TVVE using three default rendering perspectives: front, left, and top. This stage follows a similar training pipeline to RVT-2 [goyal2024rvt2], where the overall loss comprises several components. Specifically, $\mathcal{L}_{hc}$ and $\mathcal{L}_{hf}$ denote the cross-entropy losses over heatmaps produced by the coarse and fine grounding modules, respectively. The ground-truth heatmaps are generated using a truncated Gaussian distribution centered at the 2D projection of the ground-truth 3D location, as described in RVT [goyal2023rvt]. Additionally, the loss includes the end-effector rotation loss $\mathcal{L}_{rot}$, computed as the cross-entropy loss for each of the Euler angles, as well as the gripper state loss $\mathcal{L}_{gri}$ and collision indicator loss $\mathcal{L}_{col}$, both formulated as binary classification losses. The total objective for the first stage is given by:
@@
Stage 2: In the second stage, we refine the Multi-View Exploration Policy (MVEP), using the fixed-view TVVE model trained in Stage 1 as a baseline. MVEP is optimized via the Proximal Policy Optimization (PPO) algorithm [schulman2017proximal], which typically requires interaction with a physical environment to obtain reward feedback. To mitigate the high time cost, we introduce a novel pseudo-environment interaction mechanism. The details of the mechanism are presented in the supplementary materials.
@@
where $w_{i}$ are learnable component weights and $\mathcal{N}$ is an online normalization operator. The adaptive normalization maintains the mean and variance of each component using the Welford algorithm [welford1962note] for normalization, and clips the reward to the range [10,10][-10,10] to ensure training stability.
@@
C2F-ARM-BC [james2022coarse] 20.1 9.67 24.0 24.0 4.0 20.0 20.0 0.0 8.0 72.0
PerAct [shridhar2023perceiver] 49.4 7.06 $55.2{\pm}4.7$ $89.6{\pm}4.1$ $5.6{\pm}4.1$ $70.4{\pm}2.0$ $88.0{\pm}5.7$ $2.4{\pm}3.2$ $44.8{\pm}7.8$ $92.8{\pm}3.0$
HiveFormer [guhur2023instruction] 45.0 8.22 52.0 76.0 0.0 80.0 52.0 0.0 80.0 84.0
PolarNet [chen2023polarnet] 46.0 7.06 36.0 92.0 4.0 100.0 84.0 0.0 40.0 96.0
RVT [goyal2023rvt] 62.9 5.28 $52.0{\pm}2.5$ $99.2{\pm}1.6$ $11.2{\pm}3.0$ $88.0{\pm}2.5$ $71.2{\pm}6.9$ $4.0{\pm}2.5$ $91.0{\pm}5.2$ $\mathbf{100.0{\pm}0.0}$
Act3D [gervet2023act3d] 63.2 5.56 $96.8{\pm}3.0$ $80.8{\pm}6.4$ $24.0{\pm}8.4$ $95.2{\pm}1.6$ $78.4{\pm}11.2$ $3.2{\pm}3.0$ $59.2{\pm}9.3$ $93.6{\pm}2.0$
3D Diffuser Actor [ke20253d] 81.3 3.0 $96.0{\pm}2.5$ $\mathbf{100.0{\pm}0.0}$ $65.6{\pm}4.1$ $96.8{\pm}1.6$ $89.6{\pm}4.1$ $24.0{\pm}7.6$ $93.6{\pm}4.8$ $98.4{\pm}2.0$
RVT2 [goyal2024rvt2] 81.4 2.89 $\mathbf{100.0{\pm}0.0}$ $99.0{\pm}1.7$ $40.0{\pm}0.0$ $\mathbf{99.0{\pm}1.0}$ $74.0{\pm}11.8$ $38.0{\pm}4.5$ $\mathbf{95.0{\pm}3.3}$ $\mathbf{100.0{\pm}0.0}$
ARP [zhang2025autoregressive] 81.6 2.83 97.6 88.0 53.2 $96.0$ $\mathbf{90.4}$ $48.0$ $92.0$ $\mathbf{100.0}$
@@
C2F-ARM-BC [james2022coarse] 0.0 4.0 12.0 8.0 16.0 8.0 0.0 0.0 0.0 68.0
PerAct [shridhar2023perceiver] $28.0{\pm}4.4$ $51.2{\pm}4.7$ $84.0{\pm}3.6$ $17.6{\pm}2.0$ $74.0{\pm}13.0$ $16.8{\pm}4.7$ $26.4{\pm}3.2$ $2.4{\pm}2.0$ $52.0{\pm}0.0$ $88.0{\pm}4.4$
HiveFormer [guhur2023instruction] 32.0 68.0 76.0 8.0 64.0 12.0 4.0 0.0 28.0 80.0
PolarNet [chen2023polarnet] 12.0 32.0 84.0 44.0 56.0 12.0 8.0 8.0 52.0 80.0
RVT [goyal2023rvt] $49.6{\pm}3.2$ $88.0{\pm}5.7$ $91.2{\pm}3.0$ $48.0{\pm}5.7$ $81.6{\pm}5.4$ $36.0{\pm}2.5$ $28.8{\pm}3.9$ $26.4{\pm}8.2$ $72.0{\pm}0.0$ $93.6{\pm}4.1$
Act3D [gervet2023act3d] $67.2{\pm}3.0$ $91.2{\pm}6.9$ $95.2{\pm}4.0$ $32.8{\pm}6.9$ $96.0{\pm}2.5$ $29.6{\pm}3.2$ $4.0{\pm}3.6$ $9.6{\pm}6.0$ $86.4{\pm}6.5$ $94.4{\pm}2.0$
3D Diffuser Actor [ke20253d] $\mathbf{85.6{\pm}4.1}$ $96.0{\pm}3.6$ $\mathbf{97.6{\pm}2.0}$ $82.4{\pm}2.0$ $97.6{\pm}3.2$ $44.0{\pm}4.4$ $68.3{\pm}3.3$ $47.2{\pm}8.5$ $84.0{\pm}4.4$ $99.2{\pm}1.6$
RVT2 [goyal2024rvt2] $66.0{\pm}4.5$ $96.0{\pm}0.0$ $96.0{\pm}2.8$ $\mathbf{88.0{\pm}4.9}$ $92.0{\pm}2.8$ $35.0{\pm}2.8$ $\mathbf{80.0{\pm}2.8}$ $\mathbf{69.0{\pm}5.9}$ $\mathbf{100.0{\pm}0.0}$ $99.0{\pm}1.7$
ARP [zhang2025autoregressive] 68.0 $99.2$ 94.4 $85.6$ $98.4$ $35.2$ 55.2 $76.8$ $90.4$ $\mathbf{100.0}$
@@
GNFactor [ze2023gnfactor] 31.7 3.9 25.3 76.0 28.0 57.3 50.7 20.0 0.0 37.3 18.7 4.0
Act3D [gervet2023act3d] 65.3 2.7 $52.0_{\pm 5.7}$ $84.0_{\pm 8.6}$ $80.0_{\pm 9.8}$ $66.7_{\pm 1.9}$ $64.0_{\pm 5.7}$ $\mathbf{100.0_{\pm 0.0}}$ $54.7_{\pm 3.8}$ $86.7_{\pm 1.9}$ $64.0_{\pm 1.9}$ $0.0_{\pm 0.0}$
3D Diffuser Actor [ke20253d] 78.4 1.7 82.7±1.9 $\mathbf{89.3_{\pm 7.5}}$ $\mathbf{94.7_{\pm 1.9}}$ 88.0±5.7 80.0±8.6 92.0±0.0 77.3±3.8 $\mathbf{98.7_{\pm 1.9}}$ 69.3±5.0 12.0±3.7
TVVE(Ours) $\mathbf{83.2_{\pm 1.7}}$ 1.6 $\mathbf{100.0_{\pm 0.0}}$ $84.0_{\pm 5.7}$ $94.0_{\pm 2.8}$ $\mathbf{92.0_{\pm 0.0}}$ $\mathbf{92.0_{\pm 5.7}}$ $90.0_{\pm 2.8}$ $\mathbf{98.0_{\pm 2.8}}$ $50.0_{\pm 2.8}$ $\mathbf{82.0_{\pm 2.8}}$ $\mathbf{50.0_{\pm 8.5}}$

### v5

@@ Abstract
Recent vision-language-action (VLA) models for multi-task robot manipulation often rely on fixed camera setups and shared visual encoders, which limit their performance under occlusions and during cross-task transfer. To address these challenges, we propose Task-aware Virtual View Exploration (TVVE), a framework that learns to select task-relevant virtual camera viewpoints and dynamically re-render observations from a reconstructed scene representation using the selected viewpoints. To enable efficient view selection, we train an exploration policy in a pseudo-environment. In addition, we introduce a Task-aware Mixture-of-Experts (TaskMoE) visual encoder that routes visual features to task-specialized experts, mitigating interference in multi-task learning. To evaluate robustness under distribution shifts, we construct RLBench-OG, an out-of-distribution benchmark with visual perturbations and camera pose variations. Experiments on RLBench and RLBench-OG demonstrate that TVVE achieves higher success rates than strong baselines, while real-robot experiments further confirm its robustness to visual disturbances and unseen instructions. Code and visualizations are available at: TAVP.
footnotetext: Equal Contribution     Corresponding author
Accepted by CVPR 2026
@@ 1 Introduction
A general-purpose robotic system is expected to accurately perform numerous manipulation tasks [32]. With the development of vision-language-action (VLA) models [27, 3, 2, 29, 7, 4, 28, 5, 6, 31, 47, 30] and transformer-based architectures [57, 3, 2, 20, 9], many end-to-end robotic manipulation frameworks (e.g., OpenVLA[27], $\pi_{0}$[1]) have achieved promising results on complex and fine-grained tasks. Typically, these models rely on observations captured from a single view [27, 9, 44] or a few fixed viewpoints [24, 45] to guide action prediction. While this may suffice in simple environments with only a few objects, it becomes problematic in cluttered or dynamic scenes—particularly when objects or the end-effector are repositioned during task execution.
@@
Beyond acquiring complete and informative visual features, a scalable and task-sensitive visual encoder is also critical for accurate perception in multi-task settings. Although prior language-conditioned transformer-based methods, such as RVT [16] and RVT-2 [15], enable multi-task capabilities by attending to instruction tokens, their shared encoders still suffer from task interference, especially when handling visually and semantically diverse tasks (e.g., pick an apple vs. open a drawer). This entanglement limits the generalization and scalability of such models across a broad range of manipulation behaviors.
@@
To comprehensively evaluate the effectiveness of our TVVE in enhancing robotic spatial understanding and action generation performance, as well as its multi-task perception and decision-making capabilities, we conducted rigorous benchmark testing on 18 diverse RLBench [22] simulation tasks covering multiple tasks and scenarios. Our contributions are summarized as follows:
  • Multi-Viewpoint Exploration Policy (MVEP) : The MVEP enables robots to explore better virtual observation viewpoints (dynamic multi-view re-rendering), effectively addressing occlusion and insufficient viewpoint problems, and enhancing 3D perception capabilities.
@@ 2.1 Multi-task Learning in Manipulation
The core challenges for general-purpose robots lie in task generalization capabilities and multi-task execution in complex environments. In the past, multi-task learning [11, 52, 15, 35] has achieved remarkable progress. It primarily consists of two approaches: Modular solutions decompose skills into reusable basic units (e.g., perception, planning, and control modules) [36]. Though interpretable, they face the combinatorial explosion problem—as task complexity increases, the cost of designing interfaces and coordinating between modules grows exponentially, leading to system rigidity. For instance, ManipGen [10] consumed massive training resources to train over 1,000 task-specific experts. Conversely, end-to-end solutions employ a single dense model to directly map perception to action [47, 53, 49, 9, 44, 37, 19]. While simplifying the pipeline, significant differences in visual feature distributions and motion patterns across tasks (e.g., grasping fragile objects vs. tightening screws) cause the model to struggle with convergence during training due to parameter conflicts. The Mixture of Experts (MoE) [43] addresses the scaling bottleneck of large models through a sparse activation mechanism. SDP [50] integrated MoE into diffusion policies to address multi-task learning and task transfer. Differently, we propose a task-aware MoE (TaskMoE) to resolve feature conflicts between different robotic manipulation tasks, activating distinct task-specific experts for different tasks, thereby enhancing multi-task execution performance and generalization capabilities.
@@ 2.2 Vision-Language-Action Model
Recently, numerous VLA models [44, 47, 53, 15, 55] have been proposed. In contrast to traditional pipelines that rely on explicit pose estimation [13, 12] and path planning [39, 56], VLA models directly map linguistic instructions and visual observations to low-level actions, substantially simplifying the control stack. With the advancement of large language models (LLMs) and vision-language models (VLMs), these large models have been applied in robotics [27, 3, 2, 29, 7, 4], leveraging their powerful priors to provide world knowledge for robots, thereby enhancing model generalization and performance. However, most existing VLA models operate on a single or fixed viewpoint, limiting a robot’s ability to form comprehensive 3D spatial understanding and thereby constraining action prediction. To overcome this limitation, we introduce a dynamic multi-view re-rendering perception scheme that enhances spatial perception and scene comprehension, ultimately improving execution robustness in complex environments.
@@ 2.3 Reinforcement Learning in Robotics
Reinforcement learning (RL) has achieved significant accomplishments in robotics [46]. However, implementing RL starting from scratch typically necessitates complex reward mechanism and training paradigm design [21, 26, 34, 41, 48]. Recently, several works [18, 33] have demonstrated promising results by using RL to fine-tune pre-trained models. To optimize our MVEP (Multi-Viewpoint Exploration Policy), we designed pseudo-environment interactions and a view-aware reward function, enabling policy learning based on a pre-trained model using offline expert data. Furthermore, to adapt between the MVEP and the action policy, we introduce supervised post-training to coordinate the robot’s ability to “See and Act”, thereby enhancing its task execution performance.
@@ 3.1 Overview
Our proposed Task-aware Virtual View Exploration (TVVE) framework aims to identify better viewpoints for accurate and robust robotic manipulation, guided by task-specific visual feature extraction. The full pipeline is illustrated in Fig. 2. TVVE takes as input a language instruction, the current visual observations from RGB-D cameras, and the current gripper state. To enable the model to explore the better observations from arbitrary viewpoints, we first reconstruct a 3D point cloud of the scene using the input RGB-D images. To narrow down the search space and enable task-aligned view selection, we leverage the coarse prediction stage from RVT-2 [15] to generate an area of interest. However, unlike RVT-2, which extracts visual features for all tasks using a shared Multi-View Transformer (MVT), we integrate a Task-Aware Mixture-of-Experts (TaskMoE) module within the MVT to route instructions to specialized expert encoders. This design enables more precise and task-aligned visual feature extraction, which benefits both viewpoint selection and action prediction. Starting from the identified area of interest, we employ a Multi-Viewpoint Exploration Policy (MVEP) to search for better camera poses that maximize the visibility of the target object and the end-effector. The selected viewpoints are then re-rendered into image observations and processed using another TaskMoE-based MVT before being passed to the action prediction model. For action prediction, we also upgrade the autoregressive action policy proposed in ARP [55] with our proposed TaskMoE to attain more task-specific feature extraction and action prediction.
@@ 3.2 TaskMoE
To address the inherent heterogeneity of complex manipulation tasks in multi-task learning, where different tasks often require substantially distinct visual representations and action policies, we introduce a task-aware Mixture-of-Experts module (TaskMoE), as illustrated in Fig. 3. Compared to prior MoE-based approaches [43, 50], our TaskMoE introduces two key innovations.
First, instead of relying solely on task identifiers for expert selection, we incorporate richer instruction- and scene-related cues to guide expert routing more effectively, which is crucial for accurate multi-task robotic manipulation. Specifically, as shown in Fig. 3, we design a cross-modality module that employs a cross-attention mechanism to model the interaction between instruction and visual information. The resulting context-aware features are then fused with the task identifier via a Feature-wise Linear Modulation (FiLM) layer [38], enabling more adaptive and task-sensitive expert selection.
@@ 3.3 Multi-Viewpoint Exploration Policy
In this work, we represent each camera pose using a look-at model [54], which decouples the camera position and orientation via spherical coordinates. Therefore, the explored views are parameterized as 5-dimensional vectors $\mathbf{p}^{i}=(\theta^{i},\phi^{i},r^{i},\theta_{\text{up}}^{i},\phi_{\text{up}}^{i})\in\mathbb{R}^{5}$, where $(\theta,\phi,r)$ specify the camera center in spherical coordinates relative to the origin, $(\theta_{\text{up}},\phi_{\text{up}})$ define the orientation of the up vector, and $i\in[0,K-1]$ means the $i_{th}$ view. So, all camera poses are $\mathbf{p}=\{\mathbf{p}^{i}|i=0,...,K-1\}$. Additionally, the camera is perpetually oriented toward the coordinate origin $O(0,0,0)$.
@@ 3.4 Training Strategy
Stage 1: In the first stage, we train a fixed-view variant of TVVE using three default rendering perspectives: front, left, and top. This stage follows a similar training pipeline to RVT-2 [15], where the overall loss comprises several components. Specifically, $\mathcal{L}_{hc}$ and $\mathcal{L}_{hf}$ denote the cross-entropy losses over heatmaps produced by the coarse and fine grounding modules, respectively. The ground-truth heatmaps are generated using a truncated Gaussian distribution centered at the 2D projection of the ground-truth 3D location, as described in RVT [16]. Additionally, the loss includes the end-effector rotation loss $\mathcal{L}_{rot}$, computed as the cross-entropy loss for each of the Euler angles, as well as the gripper state loss $\mathcal{L}_{gri}$ and collision indicator loss $\mathcal{L}_{col}$, both formulated as binary classification losses. The total objective for the first stage is given by:
@@
Stage 2: In the second stage, we refine the Multi-View Exploration Policy (MVEP), using the fixed-view TVVE model trained in Stage 1 as a baseline. MVEP is optimized via the Proximal Policy Optimization (PPO) algorithm [42], which typically requires interaction with a physical environment to obtain reward feedback. To mitigate the high time cost, we introduce a novel pseudo-environment interaction mechanism. The details of the mechanism are presented in the supplementary materials.
@@
where $w_{i}$ are learnable component weights and $\mathcal{N}$ is an online normalization operator. The adaptive normalization maintains the mean and variance of each component using the Welford algorithm [51] for normalization, and clips the reward to the range [10,10][-10,10] to ensure training stability.
@@
C2F-ARM-BC [23] (CVPR) 20.1 9.67 24.0 24.0 4.0 20.0 20.0 0.0 8.0 72.0
PerAct [45] (CoRL) 49.4 7.06 $55.2{\pm}4.7$ $89.6{\pm}4.1$ $5.6{\pm}4.1$ $70.4{\pm}2.0$ $88.0{\pm}5.7$ $2.4{\pm}3.2$ $44.8{\pm}7.8$ $92.8{\pm}3.0$
HiveFormer [17] (CoRL) 45.0 8.22 52.0 76.0 0.0 80.0 52.0 0.0 80.0 84.0
PolarNet [8] (CoRL) 46.0 7.06 36.0 92.0 4.0 100.0 84.0 0.0 40.0 96.0
RVT [16] (CoRL) 62.9 5.28 $52.0{\pm}2.5$ $99.2{\pm}1.6$ $11.2{\pm}3.0$ $88.0{\pm}2.5$ $71.2{\pm}6.9$ $4.0{\pm}2.5$ $91.0{\pm}5.2$ $\mathbf{100.0{\pm}0.0}$
Act3D [14] (CoRL) 63.2 5.56 $96.8{\pm}3.0$ $80.8{\pm}6.4$ $24.0{\pm}8.4$ $95.2{\pm}1.6$ $78.4{\pm}11.2$ $3.2{\pm}3.0$ $59.2{\pm}9.3$ $93.6{\pm}2.0$
3D Diffuser Actor [25] (CoRL) 81.3 3.0 $96.0{\pm}2.5$ $\mathbf{100.0{\pm}0.0}$ $65.6{\pm}4.1$ $96.8{\pm}1.6$ $89.6{\pm}4.1$ $24.0{\pm}7.6$ $93.6{\pm}4.8$ $98.4{\pm}2.0$
RVT2 [15] (RSS) 81.4 2.89 $\mathbf{100.0{\pm}0.0}$ $99.0{\pm}1.7$ $40.0{\pm}0.0$ $\mathbf{99.0{\pm}1.0}$ $74.0{\pm}11.8$ $38.0{\pm}4.5$ $\mathbf{95.0{\pm}3.3}$ $\mathbf{100.0{\pm}0.0}$
ARP [55] (RA-L) 81.6 2.83 97.6 88.0 53.2 $96.0$ $\mathbf{90.4}$ $48.0$ $92.0$ $\mathbf{100.0}$
@@
C2F-ARM-BC [23] (CVPR) 0.0 4.0 12.0 8.0 16.0 8.0 0.0 0.0 0.0 68.0
PerAct [45] (CoRL) $28.0{\pm}4.4$ $51.2{\pm}4.7$ $84.0{\pm}3.6$ $17.6{\pm}2.0$ $74.0{\pm}13.0$ $16.8{\pm}4.7$ $26.4{\pm}3.2$ $2.4{\pm}2.0$ $52.0{\pm}0.0$ $88.0{\pm}4.4$
HiveFormer [17] (CoRL) 32.0 68.0 76.0 8.0 64.0 12.0 4.0 0.0 28.0 80.0
PolarNet [8] (CoRL) 12.0 32.0 84.0 44.0 56.0 12.0 8.0 8.0 52.0 80.0
RVT [16] (CoRL) $49.6{\pm}3.2$ $88.0{\pm}5.7$ $91.2{\pm}3.0$ $48.0{\pm}5.7$ $81.6{\pm}5.4$ $36.0{\pm}2.5$ $28.8{\pm}3.9$ $26.4{\pm}8.2$ $72.0{\pm}0.0$ $93.6{\pm}4.1$
Act3D [14] (CoRL) $67.2{\pm}3.0$ $91.2{\pm}6.9$ $95.2{\pm}4.0$ $32.8{\pm}6.9$ $96.0{\pm}2.5$ $29.6{\pm}3.2$ $4.0{\pm}3.6$ $9.6{\pm}6.0$ $86.4{\pm}6.5$ $94.4{\pm}2.0$
3D Diffuser Actor [25] (CoRL) $\mathbf{85.6{\pm}4.1}$ $96.0{\pm}3.6$ $\mathbf{97.6{\pm}2.0}$ $82.4{\pm}2.0$ $97.6{\pm}3.2$ $44.0{\pm}4.4$ $68.3{\pm}3.3$ $47.2{\pm}8.5$ $84.0{\pm}4.4$ $99.2{\pm}1.6$
RVT2 [15] (RSS) $66.0{\pm}4.5$ $96.0{\pm}0.0$ $96.0{\pm}2.8$ $\mathbf{88.0{\pm}4.9}$ $92.0{\pm}2.8$ $35.0{\pm}2.8$ $\mathbf{80.0{\pm}2.8}$ $\mathbf{69.0{\pm}5.9}$ $\mathbf{100.0{\pm}0.0}$ $99.0{\pm}1.7$
ARP [55] (RA-L) 68.0 $99.2$ 94.4 $85.6$ $98.4$ $35.2$ 55.2 $76.8$ $90.4$ $\mathbf{100.0}$
@@
GNFactor [52] (CoRL) 31.7 4.90 25.3 76.0 28.0 57.3 50.7 20.0 0.0 37.3 18.7 4.0
Act3D [14] (CoRL) 65.3 3.35 $52.0_{\pm 5.7}$ $84.0_{\pm 8.6}$ $80.0_{\pm 9.8}$ $66.7_{\pm 1.9}$ $64.0_{\pm 5.7}$ $\mathbf{100.0_{\pm 0.0}}$ $54.7_{\pm 3.8}$ $86.7_{\pm 1.9}$ $64.0_{\pm 1.9}$ $0.0_{\pm 0.0}$
3D Diffuser Actor [25] (CoRL) 78.4 2.20 82.7±1.9 $\mathbf{89.3_{\pm 7.5}}$ $\mathbf{94.7_{\pm 1.9}}$ 88.0±5.7 80.0±8.6 92.0±0.0 77.3±3.8 $\mathbf{98.7_{\pm 1.9}}$ 69.3±5.0 12.0±3.7
RVT2 [15] (RSS) $82.7_{\pm 0.7}$ 2.40 $96.0_{\pm 0.0}$ $78.7_{\pm 3.8}$ $73.3_{\pm 5.0}$ $\mathbf{93.3_{\pm 1.9}}$ $\mathbf{93.3_{\pm 3.8}}$ $89.3_{\pm 1.9}$ $73.3_{\pm 1.9}$ $76.0_{\pm 0.0}$ $\mathbf{96.0_{\pm 0.0}}$ $\mathbf{57.3_{\pm 3.8}}$
TVVE (Ours) $\mathbf{83.2_{\pm 1.7}}$ 2.15 $\mathbf{100.0_{\pm 0.0}}$ $84.0_{\pm 5.7}$ $94.0_{\pm 2.8}$ $92.0_{\pm 0.0}$ $92.0_{\pm 5.7}$ $90.0_{\pm 2.8}$ $\mathbf{98.0_{\pm 2.8}}$ $50.0_{\pm 2.8}$ $82.0_{\pm 2.8}$ $50.0_{\pm 8.5}$